OpenStudy (anonymous):

IF P + 77 = 101, THEN P =

Nnesha (nnesha):

how would you cancel out 77 from left side ?

OpenStudy (anonymous):

Basically what you want to do is cancel out 77 or in simpler terms get the 77 out to make the equation more easier to view. So p +77=101 so do the opposite of +77 which is -77. Subtract 77 from 101 to get 24. Now we can check out answer does 24+77=101??

OpenStudy (mindblast3r):

101-77 = P

OpenStudy (anonymous):

 (A visual to add on to it)